2.12 x 1014 protons and 3.55 đť‘Ą 1016 electrons; the charge on this object is ____


Coulombs.


Show your solution and answer to acquire the charge on the object in


Coulombs.



4.6325 x 1015 protons and 4.6325 x 1015 electrons; the charge on this object is


____ Coulombs.


Show your solution and answer to acquire the charge on the object in


Coulombs.

1
Expert's answer
2022-05-20T04:53:07-0400

(a)

"q_{net}=1.6*10^{-19}(2.12*10^{14}-3.55*10^{16})\\
=-5.65*10^{-3}\:\rm C"

(b)

"q_{net}=1.6*10^{-19}(4.6325*10^{15}-4.6325*10^{15})\\
=0\:\rm C"